OpenStudy (anonymous):

help please 5{-2+3[4-2(3+5)]} simplify

OpenStudy (anonymous):

please show steps if you can i would really appreciate it

OpenStudy (anonymous):

start working with the inner most bracketed term. slowly work your way out. 5[-2 + 3[4-2(8)]] = 5[-2+3[4-16]] = 5[-2+3(-12)] = 5[-2 -36] = 5(-38) = -190 hope that helped :)

OpenStudy (anonymous):

thanks a bunch xD

OpenStudy (amistre64):

5{-2+3[4-2(3+5)]} -10+15[4-2(3+5)] -10+60-30(3+5) 50 -90-150 -90-100 -190 maybe?

OpenStudy (anonymous):

your equestion is wrong it has to be like 5[-2+3{4-2(3+5)}]

OpenStudy (amistre64):

lol ... bad equation!, bad!

OpenStudy (anonymous):

that's how it's written in my book so bad editors bad!

OpenStudy (amistre64):

.... there is no order to use of brackets. Just as long as you match pairs your fine

OpenStudy (anonymous):

lol , i hate equations like that since i start school :P

OpenStudy (anonymous):

they can get pretty confusing

OpenStudy (amistre64):

5(-2 +3(4 -2(3+5))), is just as acceptable, the different patterns tho help to see the groupings better, thats all

OpenStudy (anonymous):

well i think to put some spaces between the brackets will make you not to get confused :D

OpenStudy (anonymous):

true

OpenStudy (anonymous):

5[-2+3{4-2(3+5)}] Or 5[-2+3{4-2*8}] Or 5[-2+3{4-16}] Or 5[-2 +3* -12] Or 5[-2-36] Or 5*-38 Or -190

OpenStudy (amistre64):

d<1+c<2+b<3+a>>> lol

OpenStudy (anonymous):

lol 4real

OpenStudy (anonymous):

i go solve some integral its so confusing :P

OpenStudy (anonymous):

(it = that question here :D )